From nch@roe.ac.uk Fri May 14 18:03:17 2010 Date: Fri, 14 May 2010 15:30:27 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 14/5/10 Minutes of WFAU VDFS Science Archive meeting: May 14th 2010 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, RSC, RPB, MAR, NJC Apologies: JPE, AL, MSH, ETWS, KTN, RGM DoNM: 10am, Friday May 21st 2010 in the VISTA Hut Actions discharged this week: ----------------------------- ACTION: NCH to email PS PIs proposing a formal (and efficient!) method for user-specified QC deprecations in the VSA. Discharged; no substantive responses so far... ACTION: RPB to ask RGMcM if he can guarantee that the VHS survey team consists of ESO astronomers only Discharged; RGMcM assures he won't give out access credentials to non-ESO astronomers, but has requested crossmatch with both world DR4 *and* proprietary DR7 anyway. RPB says "no problem". Actions partly discharged but continuing: ----------------------------------------- None this week. Actions carried forward from 07/05/10 meeting: ---------------------------------------------- ACTION: NCH to fill out a Q2 2010 plan and continue with the progress monitoring for the time being. Continues ACTION: RPB/KTN to order several crates of LTO-4 tapes. Continues ACTION: ALL to consider who goes to Garching in September for the VST PS readiness meeting. Continues Specific points and new actions: -------------------------------- Project management: RGM's communication with JPE regarding co-ordination of archive-end things with PSPIs received a positive response, so NJC will convene a regular (every couple of months) Skype teleconf with the PIs, JPE and MAR to thrash out any issues of importance like QC co-ordination, release co-ordination etc. ACTION: NJC to Doodle poll for a kick-off PSPI co-ordination teleconf WFCAM & VISTA updates: Nothing of note this week. Comments and issues arising from CASU minutes: No new minutes this week. Networking: RPB noted that VISTA data is all copied and metadata ingested up to end March v0.8 (including the P85 survey-proper data); WFCAM transfers are all up-to-date with 10A processing. JAC have indicated the appearance of new focus metadata keyword FOC_TOFF in WFCAM headers - this needs adding into the WSA multiframe schema. WSA/VSA Operations: NCH noted that given the entire ops team will have FOC_TOFF on holiday next week, those of us that remain need to try to keep the ball rolling on archive curation activities. After a long discussion, RPB (and MAR for QC/UI admin) agreed to email around some hand-holding instructions so that preparation of catalogue release DBs for VISTA and web server functions can continue. NCH will keep an eye on things (note to self: curation logs appear on /disk33/vsa/logs) and will be the holder of the scos password for the duration. The agreed priority is VVV, VIKING, VMC, then fresh releases of VHS and VIDEO incorporating the latest data and changes as requested by PIs. UltraVista and SV are shelved in lieu of v1.0 processing unless specific requests come in from the respective PIs. RSC volunteered to slot in some much-delayed WSA non-survey deblend fixing in the event of any spare CPU cycles. RPB/MAR noted that some cross-neighbour funtionality between VSA DBs and external catalogues like SSA/SDSS has been non-operational up to now, but will be fixed today given that things have stabilised sufficiently. Hardware and Systems: NJC noted that the main bottleneck in curation seems to be availability of processing power for the deep-stacked VISTA surveys. RPB noted that a new beefy curation server is in the racks awaiting set-up, and that this should be MSH's top priority when he returns from his riotous holiday in Greece later next week. NCH noted that we should keep the situation under review ACTION: ALL to review new hardware requirements/procurement over the next couple of weeks Software: RSC noted that the new release branch creation in SVN is on hold while NJC investigates a little bug in autocurate, and that a small mod to the CU19 release script to auto-generate DB statistics is also on hold since MSH indicated earlier in the week that he'd found some cases where query execution plans were actually slower in the presence of stats. NCH wondered if this was the old SQL 2008 parallelism issue coming back to haunt us. In any case, we await MSH's return before making a final decision. Survey Data Release: MAR noted that a new VIDEO release had been made in the week, including pawprint deep stack, band-merged and variability products. Non-survey Data Release: MAR noted that archive flat-file access for WFCAM non-survey PIs is all up-to-date with processing in semester 10A. Astrogrid deployment & Data Analysis services: Nothing of note this week. Miscellaneous: Nothing else this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================